# Prox-DBRO-VR: A Unified Analysis on Decentralized Byzantine-Resilient Composite Stochastic Optimization with Variance Reduction and Non-Asymptotic Convergence Rates

### v6

@@ Abstract
Decentralized Byzantine-resilient stochastic gradient algorithms resolve efficiently large-scale optimization problems in adverse conditions, such as malfunctioning agents, software bugs, and cyber attacks. This paper aims to handle a class of general composite optimization problems over multi-agent cyber-physical systems (CPSs), with the existence of an unknown number of Byzantine agents. Based on the proximal mapping method, variance-reduced (VR) techniques, and a norm-penalized approximation strategy, we propose a decentralized Byzantine-resilient and proximal-gradient algorithmic framework, dubbed Prox-DBRO-VR, which achieves an optimization and control goal using only local computations and communications. To reduce asymptotically the variance generated by evaluating the noisy stochastic gradients, we incorporate two localized variance-reduced techniques (SAGA and LSVRG) into Prox-DBRO-VR to design Prox-DBRO-SAGA and Prox-DBRO-LSVRG. By analyzing the contraction relationships among the gradient-learning error, robust consensus condition, and optimal gap, theoretical results demonstrate that both Prox-DBRO-SAGA and Prox-DBRO-LSVRG, with a well-designed constant (resp., decaying) step-size, converge linearly (resp., sub-linearly) inside an error ball around the optimal solution to the optimization problem under standard assumptions. The trade-off between convergence accuracy and the number of Byzantine agents in both linear and sub-linear cases is characterized. In simulation, the effectiveness and practicability of the proposed algorithms are manifested via resolving a decentralized sparse machine-learning problem over multi-agent CPSs under various Byzantine attacks.
@@
Decentralized stochastic optimization, system security, cyber attacks, Byzantine-resilient algorithms, composite optimization problems, variance-reduced stochastic gradients.
@@ I-A Literature Review
Decentralized optimization has seen extensive research and notable progress in the field of machine learning [1, 2, 3], smart grids [4], cooperative control [5], and uncooperative games [6]. Decentralized optimization algorithms have the advantages of high-efficiency for massive-scale optimization problems, good scalability over large-scale intelligent systems, and a lower cost in short-distance communications.
@@
With the rapid expansion of multi-agent CPSs, there are unavoidable security issues in the process of optimization and control, such as poisoning data, software bugs, malfunctioning devices, and cyber attacks [7, 8]. All these issues in the course of multi-agent optimization and control are generalized as a node-level problem model, namely Byzantine problems [9], while the malfunctioning or compromised agents are called Byzantine agents. Byzantine agents are able to stop the existing notable decentralized optimization algorithms [10, 11, 2, 5, 12, 13, 14, 6, 15, 4, 16, 1, 17] from achieving the optimal solution to the optimization problem [18], or even cause disagreement and divergence [19]. For example, if an honest agent is attacked and controlled by adversaries, the attacker can manipulate the agent to send misleadingly falsified information to its different reliable neighboring agents at each iteration. This can easily deter reliable agents from achieving convergence, and even the consensus is impeded if one Byzantine agent chooses to send different misleading messages to different reliable neighbors. Therefore, researchers have been concentrating on designing decentralized resilient algorithms [20, 21, 22, 23, 24, 25, 26, 27, 28] to alleviate or counteract the negative impact caused by Byzantine agents. In fact, there are various approaches to guarantee decentralized Byzantine resilience. One popular line is to combine various screening or filtration techniques with decentralized algorithms. To name a few, ByRDiE [29] requires that each agent at every iteration discards a subset of the largest and smallest messages in the received information, which is followed by a coordinate gradient descent step. One imperfection of ByRDiE is its expensive computational overhead and low efficiency in dealing with large-scale optimization problems due to the implementation of one-coordinate-at-one-iteration update. Hence, BRIDGE [23] combines respectively four screening techniques including coordinate-wise trimmed-mean, coordinate-wise median, Krum function, and a combination of Krum and coordinate-wise trimmed mean, with decentralized gradient descent (DGD) [30] to devise a Byzantine-resilient algorithmic framework. However, these four screening mechanisms either suffer from a high computational complexity or impose extra restrictions on the number of neighbors and the network topology. Follow-up literature [31] incorporates a self-centered clipping technique (adapted from the centered clipping [32]) into DGD to not only realize Byzantine resilience, but resolve a category of general non-convex objective functions. However, the decentralized algorithm proposed in [31] assumes that each agent has the prior knowledge of global parameters, for instance, the subset of Byzantine agents. This may be impractical in real large-scale CPSs, since the information exchange is only available in a decentralized manner. Another literature [18] designs a two-stage technique to filter out the Byzantine attacks, which can work over multi-agent CPSs with an arbitrary quantity of Byzantine agents while any clairvoyant knowledge of the identities of Byzantine agents is not required. Recent work [26] systematically analyzes the challenges on two critical points, i.e., doubly-stochastic weight matrix and consensus, in the development of decentralized Byzantine-resilient methods. Based on the analysis, a screening-based robust aggregation rule, dubbed (IOS), is designed in [26], which achieves Byzantine resilience and a controllable convergence error relied on the assumptions of bounded inner (node-level noisy stochastic gradients) and outer (network-level aggregated gradients) variations.
@@
All of the aforementioned decentralized Byzantine algorithms [29, 23, 31, 18, 26] achieve Byzantine resilience via adopting various screening or filtering techniques. Nevertheless, the screening or filtration-based methods may not only impose restrictions on the minimum number of neighboring agents, but incur at least an additional computational cost of $\mathcal{O}\left({n\left|{{\mathcal{N}_{i}}}\right|}\right)$ ($n$ and ${\mathcal{N}_{i}}$ denote the dimension of single data sample and the number of neighboring agents of the reliable agent $i$, respectively) to each reliable agent $i$ at each iteration (see [23, TABLE II]). This could be prohibitively expensive if the optimization problem is high-dimensional (machine learning or deep learning tasks) or the multi-agent CPSs are large-scale. One possible solution to avoid introducing extra computational costs for decentralized Byzantine-resilient optimization is developed by RSA [28], which combines an $a$-norm-penalized ($a\geq 1$) approximation method with the stochastic sub-gradient descent method to realize robust aggregation in a distributed fashion. [27] is a decentralized extension of RSA [28]. Via integrating with a noise-shuffle strategy, [33] extends [28] to distributed federated learning, which enhances users’ differential privacy. However, both RSA [28] and the literature [27, 33] establish only sub-linear convergence rates of the proposed algorithms, which are rather slow and have huge potential to be accelerated. A recent decentralized Byzantine-resilient algorithm DECEMBER [34] accelerates the convergence rate via incorporating two variance-reduced techniques. Although DECEMBER realizes simultaneously Byzantine resilience and linear convergence, it is still confined to resolving optimization problems with only a single smooth objective function.
@@ I-B Motivations
All aforementioned decentralized Byzantine-resilient methods [29, 28, 27, 33, 23, 31, 18, 26, 34] are not available to handling optimization problems with the existence of any non-smooth objective, which is indispensable in many practical applications, such as sparse machine learning [35, 36], model predictive control [5], and energy resource coordination [12]. On the other hand, despite that there are various decentralized algorithms [37, 35, 38, 36, 12, 3] providing different insights and tactics to resolve the composite optimization problem, they all fail to consider any possible security issues over CPSs, which renders the reliable agents under the algorithmic framework of [35, 38, 36, 12] easily misled by Byzantine failures and vulnerable to Byzantine attacks. Therefore, to close this gap, this paper focuses on studying a category of composite optimization problems in the presence of Byzantine agents, where the local objective function associated with each agent consists of both smooth and non-smooth parts. In a nutshell, the study on decentralized Byzantine-resilient composite stochastic optimization is non-trivial, which features the main motivation of this paper.
@@ I-C Contributions
    This paper designs a decentralized Byzantine-resilient and proximal-gradient algorithmic framework, dubbed Prox-DBRO-VR, to resolve a class of composite (smooth + non-smooth) optimization problems over multi-agent CPSs under the worst case of Byzantine attacks. The worst case indicates that the number of Byzantine agents is unlimited and any information relevant to their identities is not necessarily known by reliable agents. To reduce the expensive per-iteration computational cost in decentralized Byzantine-resilient batch-gradient methods [29, 23, 39, 40, 20], we incorporate the localized versions of two variance-reduced techniques SAGA [41] and LSVRG [42], into Prox-DBRO-VR, to propose two decentralized Byzantine-resilient stochastic gradient algorithms, namely Prox-DBRO-SAGA and Prox-DBRO-LSVRG. Owing to the employment of the variance-reduced techniques, Prox-DBRO-SAGA and Prox-DBRO-LSVRG reduce asymptotically the variance incurred by the local stochastic gradients, which also eliminates the bounded-variance assumption required by [43, 28, 27, 14, 26].
@@
    In contrast to recent works [39, 23, 25, 26, 24, 34, 40, 20], this papers considers a more general composite optimization problem model with local non-smooth objective functions. This superiority also brings challenges in theoretical analysis, which has been addressed by exploring the contraction properties of proximal operators over decentralized Byzantine-resilient optimization. Different with [28, 27, 33, 34], a unified convergence analysis is conducted in this paper to obtain more intuitive and complete theoretical results, including both a sub-linear convergence rate (with a smaller convergence error) and a linear convergence rate (with a larger convergence error), which provides in-depth knowledge of the trade-off between convergence speed and convergence accuracy. Furthermore, the uncoordinated parameter setup is considered in algorithm development, i.e., uncoordinated penalty parameters of Prox-DBRO-VR and uncoordinated triggered probabilities of Prox-DBRO-LSVRG, which not only allows each agent to decide independently their own parameters, but contributes to attaining a more complete convergence result.
@@
    The screening or filtration-based Byzantine-resilient methods, such as [19, 23, 29, 39, 18, 25, 26, 24], may impose a rigorous assumption on the number of neighboring agents of the reliable agents or the network topology, which may be impractical in the large-scale CPSs with a complex system structure. To eliminate this requirement, Prox-DBRO-SAGA and Prox-DBRO-LSVRG adopt a generalized penalized-norm approximation method to realize Byzantine resilience. Moreover, both Prox-DBRO-SAGA and Prox-DBRO-LSVRG achieve Byzantine resilience without incurring any additional costs in contrast to the decentralized Byzantine-resilient methods [19, 23, 29, 25, 18, 26, 24] that result in either prohibitively expensive screening or observation costs, especially when facing large-scale or high-dimensional optimization problems. Besides, the theoretical analysis of Prox-DBRO-SAGA and Prox-DBRO-LSVRG only requires one potential connected network among reliable agents, which is more relaxed than the observation-based methods [40, 20] relying on trust observations and assumptions of sufficiently connected networks and bounded malicious information.
@@ II-A Basic Notation
Throughout the paper, we assume all vectors are column vectors if there is no other specified.
@@
${\mathbb{R}}$, ${{\mathbb{R}}^{n}}$, ${{\mathbb{R}}^{m\times n}}$
the set of real numbers, $n$-dimensional column real vectors, $m\times n$ real matrices, respectively
${I_{n}}$
the $n\times n$ identity matrix
${0_{n}}$
an $n$-dimensional column vector with all-zero elements
$1_{m}$
an $m$-dimensional column vector with all-one elements
${\cdot^{\top}}$
transpose of any matrices or vectors
${\rm{diag}}\left\{\nu\right\}$
a diagonal matrix with all the elements of vector $\nu\in{\mathbb{R}^{n}}$ laying on its main diagonal
$X\leq Y$
each element in $Y-X$ is nonnegative, where $X$ and $Y$ are two vectors or matrices with same dimensions
$\tilde{x}\otimes\tilde{y}$
the Kronecker product of vectors $\tilde{x}$ and $\tilde{y}$
@@
the operator to represent the absolute value of a constant or the cardinality of a set
$\left\|\nu\right\|_{a}$
either the $a$-norm of $\nu\in{\mathbb{R}^{n}}$ equivalent to ${\left({\sum\nolimits_{i=1}^{n}{{{\left|{{\nu_{i}}}\right|}^{a}}}}\right)^{\frac{1}{a}}}$, $a\geq 1$, or its induced matrix norm.
${\lambda_{\min}}\left(X\right)$
the minimum nonzero singular value of any matrix $X$
${\lambda_{\max}}\left(X\right)$
the maximum singular value of any matrix $X$
TABLE I: Basic notations.
For arbitrary three vectors $\tilde{x},\tilde{y},\tilde{z}\in\mathbb{R}^{n}$, a positive scalar $a$ and a closed, proper, convex function, $g:{\mathbb{R}^{n}}\to\mathbb{R}$, the proximal operator is defined as: ${\mathbf{prox}}_{{a,g}}\left\{\tilde{x}\right\}=\arg{\min_{\tilde{y}\in\mathbb{R}^{n}}}\left\{{g\left(\tilde{y}\right)+\frac{1}{{2a}}{{\left\|{\tilde{y}-\tilde{x}}\right\|}_{2}^{2}}}\right\}$; let $\partial g\left(\tilde{x}\right)$ denote the sub-differential of the proper, closed and convex function $g:{\mathbb{R}^{n}}\to\mathbb{R}$ at $\tilde{x}$, such that
$\partial g\left(\tilde{x}\right)=\left\{{\tilde{y}|\forall\tilde{z}\in{\mathbb{R}^{n}},\;g\left(\tilde{x}\right)+\left\langle{\tilde{y},\tilde{z}-\tilde{x}}\right\rangle\leq g\left(\tilde{z}\right)}\right\};$
let ${{\partial_{\tilde{x}}}g\left(\tilde{x}\right)}$ denote the sub-gradient of non-smooth convex function $g$ at $\tilde{x}$. The remaining basic notations of this paper are summarized in Table I.

### v12

# Prox-DBRO-VR: A Unified Analysis on Byzantine-Resilient Decentralized Stochastic Composite Optimization with Variance Reduction

@@ Abstract
Decentralized stochastic gradient algorithms efficiently solve large-scale finite-sum optimization problems when all agents in the network are reliable. However, most of these algorithms are not resilient to adverse conditions, such as malfunctioning agents, software bugs, and cyber attacks. This paper aims to handle a class of general composite optimization problems over multi-agent systems (MASs) in the presence of an unknown number of Byzantine agents. Building on a resilient aggregation mechanism and the proximal-gradient mapping method, a Byzantine-resilient decentralized stochastic proximal-gradient algorithmic framework is proposed, dubbed Prox-DBRO-VR, which achieves an optimization and control goal using only local computations and communications. To asymptotically reduce the noise variance arising from local gradient estimation and accelerate the convergence, we incorporate two localized variance-reduced (VR) techniques (SAGA and LSVRG) into Prox-DBRO-VR to design Prox-DBRO-SAGA and Prox-DBRO-LSVRG. By analyzing the contraction relationships among the gradient-learning error, resilient consensus condition, and convergence error in a unified theoretical framework, it is proved that both Prox-DBRO-SAGA and Prox-DBRO-LSVRG, with a well-designed constant (resp., decaying) step-size, converge linearly (resp., sub-linearly) inside an error ball around the optimal solution to the original problem under standard assumptions. A trade-off between convergence accuracy and Byzantine resilience in both linear and sub-linear cases is also characterized. In numerical experiments, the effectiveness and practicability of the proposed algorithms are manifested via resolving a decentralized sparse machine-learning problem under various Byzantine attacks.
@@
Decentralized stochastic optimization, system security, Byzantine-resilient algorithms, composite optimization, variance reduction.
@@ I-A Literature Review
Decentralized optimization has garnered significant attention and achieved substantial algorithmic breakthroughs in the fields of machine learning [1, 2, 3], smart grids [4], cooperative control [5], and uncooperative games [6]. Decentralized algorithms have advantages of high-efficiency in massive-scale optimization problems, good scalability over geographically distributed MASs, and a lower communication burden for the master/central agent in a parameter-server distributed structure.
@@
The increasing deployment of MASs in aforementioned mission-critical applications, there are unavoidable security issues in the course of optimization and control, such as poisoning data, software bugs, malfunctioning devices, cyber attacks, and privacy leakage. Many endeavors have been devoted to develop various privacy-preserving methods in decentralized optimization, for instance, differential privacy [7, 8]. Nevertheless, privacy issues are beyond the scope of this paper, as they neither directly cause compromised agents nor violate the prescribed update rules of algorithms. Other issues, however, may lead to node-level failures during multi-agent optimization and control, which is known as Byzantine problems [9], where the malfunctioning or compromised agents are referred to as Byzantine agents. Byzantine agents colluding with each other are able to impede some notable Byzantine-free (non-resilient) decentralized optimization algorithms [10, 11, 2, 5, 12, 13, 14, 6, 15, 4, 1, 16] from achieving convergence, or even cause disagreement and divergence [17, 18]. For example, if a reliable agent is attacked and controlled by adversaries, the attacker can manipulate the agent to send misleadingly falsified information to its different reliable neighboring agents at each iteration. This can significantly disrupt the normal update of reliable agents, and even stop them from reaching consensus if a Byzantine agent deliberately sends misleading messages to their different reliable neighbors [19]. Therefore, researchers have been concentrating on designing resilient decentralized algorithms [20, 21, 22, 23, 24, 25, 26, 27, 28] to alleviate or counteract the negative impact caused by Byzantine agents. In fact, there are various approaches to secure resilience in a decentralized manner. One popular line is to combine various screening or filtration techniques with decentralized optimization algorithms. To name a few, a pioneering work [17] achieves Byzantine resilience via adopting trimmed mean (TM) aggregation, which discards a subset of the largest and smallest messages in the aggregation step. However, [17] focuses on a class of scalar-valued optimization problems for decentralized optimization. A follow-up method ByRDiE [29] combines decentralized coordinate gradient descent with TM aggregation to tackle the vector-valued optimization problem in decentralized learning. One imperfection of ByRDiE stems from low efficiency in handling large-scale finite-sum optimization problems due to the implementation of one-coordinate-at-one-iteration update. Hence, BRIDGE [23] combines respectively four screening techniques including coordinate-wise trimmed-mean, coordinate-wise median, Krum function, and a combination of Krum and coordinate-wise trimmed mean, with decentralized gradient descent (DGD) [30]. However, these four screening mechanisms either incur prohibitive computational overhead or impose extra restrictions on the number of neighbors or the network topology. To address these issues, building upon a centered clipping (CC) technique [31], He et al. [32] propose self-centered clipping (SCC) for decentralized implementations, which not only ensures Byzantine resilience but also provides a complete solution framework for decentralized non-convex optimization problems. One imperfection of SCC [31] is that the theoretical choice of localized clipping parameters requires knowledge of information that is assumed to be inaccessible. This limitation mandates manual parameter calibration in the practical deployment of SCC, thereby compromising its autonomy. Another work [33] designs a two-stage filtering technique to mitigate Byzantine threats, which is independent of any clairvoyant knowledge about Byzantine agents. Recent work [26] systematically analyzes the relationship between two critical points, i.e., doubly-stochastic weight matrix and consensus, in the development of Byzantine-resilient decentralized optimization algorithms. On top of that, an iterative screening-based resilient aggregation rule, dubbed (IOS), is designed in [26], which achieves Byzantine resilience and a controllable convergence error relied on the assumptions of bounded inner (node-level noise-based stochastic gradients) and outer (network-level aggregated gradients) variations.
@@
Decentralized optimization algorithms [29, 33, 23, 26, 25] achieve Byzantine resilience via adopting various screening or filtering techniques. Nevertheless, the screening- or filtration-based methods may either impose a restriction on the number of neighboring agents or introduce additional computational complexity of $\mathcal{O}\left(mn\right)$ ($\mathcal{O}$, $m$, and $n$ denote an upper bound on the growth rate of the complexity, the total number of agents including both reliable and Byzantine agents in the network, and the dimension of decision variables, respectively), at each iteration (see [23, TABLE II]). This could be prohibitively expensive when the MAS is large-scale or the decision variable is high-dimensional. By integrating a resilient aggregation mechanism based on norm-penalized approximation [36], RSA [28] is developed for distributed federated-learning problems in the presence of Byzantine agents, which achieves Byzantine resilience without introducing prohibitive computational overhead. [27] is a decentralized extension of RSA [28]. Via introducing a noise-shuffle strategy, [34] enables [28] to be differentially private in distributed federated-learning tasks. However, the existing theoretical evaluation reveals that the algorithms proposed in [28, 27, 34] sacrifice either convergence accuracy or convergence speed due to the employment of noise-based stochastic gradients. This dilemma can be resolved via employing VR techniques. Therefore, a recent Byzantine-resilient decentralized stochastic optimization algorithm DECEMBER [35] accelerates the convergence via incorporating VR techniques. Although DECEMBER resolves the dilemma, it is still confined to empirical risk minimization (ERM) problems with local smooth objectives.
@@ I-B Motivations
On the one hand, Byzantine-resilient decentralized optimization algorithms [36, 28, 27, 32, 33, 23, 25, 26, 24, 35, 31, 34] are not available to handling optimization problems with a nonsmooth objective function, which is indispensable in many practical applications, such as sparse machine learning [37, 38], model predictive control [5], and signal processing [39]. On the other hand, despite the fact that there are various decentralized optimization algorithms [39, 40, 37, 38, 12, 41, 3] providing many insights to resolve the composite optimization problem, they all fail to consider any possible security issues over MASs. This renders the reliable agents under the algorithmic framework of [39, 40, 37, 41, 38, 12, 3] vulnerable to Byzantine attacks or failures. To bridge this gap, this paper studies a category of composite optimization problems in the presence of Byzantine agents, where the local objective function associated with each agent consists of both smooth and nonsmooth objectives. In a nutshell, the non-trivial study on Byzantine-resilient decentralized stochastic composite optimization deserves further investigation, which features the main motivation of this paper. The integration of VR techniques serves as a side motivation, which not only helps reduce the prohibitive per-iteration gradient-computation cost when evaluating the local batch gradients [39, 29, 17, 40, 38, 41, 23, 25, 20] but also asymptotically reduces the noise variance arising from local gradient estimation [28, 32, 33, 27, 14, 31, 26, 24]. In contrast to a recent notable algorithm (event-triggered) MW-MSR [19, 42], which proposes an alternative approach to achieving Byzantine-resilient consensus with asynchronous multi-hop communication through detection and filtration, this paper focuses on decentralized optimization problems. In such problems, the heterogeneity in local gradient evaluations can lead to greater variance among reliable agents’ states [31], which may cause false filtration when the reliable information is mistakenly identified as Byzantine information. This observation motivates the design of a screening- and filtration-free Byzantine-resilient decentralized optimization algorithm tailored to handle the worst-case scenario of Byzantine problems in decentralized optimization. The worst-case scenario indicates that the number of Byzantine agents is unknown and they can be omniscient to send any misleadingly falsified messages to their reliable neighbors.
@@ I-C Contributions
    This paper develops a Byzantine-resilient decentralized stochastic proximal-gradient algorithmic framework, dubbed Prox-DBRO-VR, to resolve a class of composite (smooth + nonsmooth) finite-sum optimization problems over MASs under the worst case. The challenge of studying the nonsmooth objective function in the presence of Byzantine agents stems from incurring additional error terms consisting of coupled Byzantine and reliable information in contrast to Byzantine-free composite optimization [39, 40, 37, 38, 12, 41, 3] and Byzantine-resilient smooth optimization [36, 28, 27, 32, 33, 23, 25, 26, 24, 35, 31, 34]. To handle these introduced error terms, we explore and seek the upper bounds on the subdifferentials related to Byzantine agents after applying the non-expansiveness of the proximal operator in the linear convergence case and bounded-gradient condition in the sub-linear convergence case, respectively.
@@
    Inspired by [37, 11], we incorporate localized versions of two VR techniques SAGA [43] and LSVRG [44], into Prox-DBRO-VR, to propose two Byzantine-resilient decentralized stochastic proximal gradient algorithms, namely Prox-DBRO-SAGA and Prox-DBRO-LSVRG, both of which reduce the per-iteration gradient-computation cost in evaluating batch gradients [39, 29, 17, 40, 38, 41, 23, 25, 20] and circumvent the bounded-variance assumption required by noise-based stochastic gradient methods [28, 32, 33, 27, 14, 31, 26, 24, 34]. The challenge of studying VR techniques in the presence of Byzantine agents lies in seeking an appropriate (Lyapunov) candidate function with respect to (w.r.t.) the gradient-learning sequences and errors, we address this challenge by exploiting the Bregman divergence, choosing a proper decaying or constant step-size, and utilizing appropriate intermediate constants (see the proofs of Theorems 2-3).
@@
    The proposed algorithmic framework Prox-DBRO-VR achieves Byzantine resilience without incurring prohibitive computational overhead in contrast to Byzantine-resilient decentralized optimization algorithms [17, 29, 23, 25, 33, 26, 24, 19] introducing additional computational complexity of $\mathcal{O}\left(mn\right)$ for screening or filtration processes. The theoretical analysis of Prox-DBRO-SAGA and Prox-DBRO-LSVRG imposes no assumption or restrictions on the number or proportion of Byzantine agents in the network, while only assumes a connected network among all reliable agents (see Corollary 1), which is less restrictive than the topology condition of many related studies, such as [17, 29, 23, 25, 26]. Under this assumption, theoretical results reveal an explicit trade-off between convergence accuracy and Byzantine resilience in both cases (see Theorems 2-3 for details), providing directions to optimize the performance of the proposed algorithms in practice.
@@ II-A Basic Notation
Throughout the paper, we assume all vectors are column vectors if there is no other specified. For arbitrary three vectors $\tilde{x},\tilde{y},\tilde{z}\in\mathbb{R}^{n}$, a positive scalar $a$ and a closed, proper, convex function $g:{\mathbb{R}^{n}}\to\mathbb{R}$, the proximal operator of $g$ is defined by ${\mathbf{prox}}_{{a,g}}\left\{\tilde{x}\right\}=\arg{\min_{\tilde{y}\in\mathbb{R}^{n}}}\left\{{g\left(\tilde{y}\right)+\frac{1}{{2a}}{{\left\|{\tilde{y}-\tilde{x}}\right\|}_{2}^{2}}}\right\}$; let $\partial g\left(\tilde{x}\right)$ denote the subdifferential of $g$ at $\tilde{x}$, such that
$\partial g\left(\tilde{x}\right)=\left\{{\tilde{y}|\forall\tilde{z}\in{\mathbb{R}^{n}},g\left(\tilde{x}\right)+\left\langle{\tilde{y},\tilde{z}-\tilde{x}}\right\rangle\leq g\left(\tilde{z}\right)}\right\},$
let ${{\partial_{\tilde{x}}}g\left(\tilde{x}\right)}$ denote one subgradient of $g$ at $\tilde{x}$, such that ${{\partial_{\tilde{x}}}g\left(\tilde{x}\right)}\in{{\partial}g\left(\tilde{x}\right)}$.
TABLE I: Basic notations.
@@
${\mathbb{R}}$, ${{\mathbb{R}}^{n}}$, ${{\mathbb{R}}^{m\times n}}$ the set of real numbers, $n$-dimensional column real vectors, $m\times n$ real matrices, respectively
${\mathbb{E}_{\xi}}\left[\cdot\right]$ the expectation with respect to a random variable $\xi$
:= the definition symbol
${I_{n}}$ the $n\times n$ identity matrix
${0_{n}}$ an $n$-dimensional column vector with all-zero elements
$1_{m}$ an $m$-dimensional column vector with all-one elements
${\cdot^{\top}}$ transpose of any matrices or vectors
$X\leq Y$ each element in $Y-X$ is nonnegative, where $X$ and $Y$ are two vectors or matrices with same dimensions
$\tilde{x}\otimes\tilde{y}$ the Kronecker product of vectors $\tilde{x}$ and $\tilde{y}$
@@
$\left\|\nu\right\|_{a}$ either the $a$-norm of $\nu\in{\mathbb{R}^{n}}$ equivalent to ${\left({\sum\nolimits_{i=1}^{n}{{{\left|{{\nu_{i}}}\right|}^{a}}}}\right)^{\frac{1}{a}}}$, $a\geq 1$, or its induced matrix norm.
${\mathbf{col}}\left\{\cdot\right\}_{i}$ the operator that vertically concatenates the input vectors indexed by $i$ into one column vector
${\lambda_{\min}}\left(X\right)$ the minimum nonzero singular value of any matrix $X$
${\lambda_{\max}}\left(X\right)$ the maximum singular value of any matrix $X$
The remaining basic notations of this paper are summarized in Table I.